Whispering Inc. loans money to John Kruk Corporation in the amount of $848,000. Whispering accepts an 8% note due in 7 years with interest payable semiannually. After 2 years (and receipt of interest for 2 years), Whispering needs money and therefore sells the note to Chicago National Bank, which demands interest on the note of 10% compounded semiannually. What is the amount Whispering will receive on the sale of the note? (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.) What is the total amount received on the sale of the note?
Answer #1
Amount received on the sale of the note 782517
Workings:
Semi-annual interest 33920 =848000*8%*6/12
X PV factor 7.72173 =(1-(1.05)^-10)/0.05
Present value of interest 261921
Principal 848000
X PV factor 0.61391 =1/1.05^10
Present value of principal 520596
Amount received on the sale of the note 782517 =261921+520596